OpenStudy (anonymous):

HELP PLEASE!! Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (mathmate):

How much does the club make selling ONE flower: cost = 100% = 0.50 profit = 40% = (40/100)*0.50=0.20 So profit for 1 flower = 0.20 profit for 5 flowers = 5*0.20 = $1.00 How many flowers do they need to sell to make $500 ?

OpenStudy (anonymous):

Ok so...